OpenStudy (anonymous):

If f (θ) = tan θ, find the differential df when θ = π/3 and dθ = 0.03.

OpenStudy (anonymous):

\[\Delta f = \tan(\theta+\Delta\theta) - \tan(\theta)\]

OpenStudy (anonymous):

tan(0.03+delta0.03)-tan(0.03)????

OpenStudy (anonymous):

\[\tan(\frac\pi3 + 0.03) - \tan(\frac\pi3)\]

OpenStudy (anonymous):

0.12

OpenStudy (anonymous):

0.1266, which rounds to 0.13 but essentially right

OpenStudy (anonymous):

okay awesome! Thanks!